find the first second and third number of
1/3x+1/7(x+1)=1/2(x+2)-2
find the smallest and larger number of
1/3(x+2)-1/5x=12
find the first second and third number of
x+1/2(x2)+1/4(x+4)=2x-28

Answer by stanbon(75887)

find the first second and third number of
1/3x+1/7(x+1)=1/2(x+2)-2
---
Multiply thru by 42 to get:
14x + 6(x+1) = 21(x+2)-84
---
20x + 6 = 21x-42
x = 48
--
x+1 = 49
---
x+2 = 50
=================
Cheers,
Stan H.
==================